# Detection of extended X-ray emission surrounding PSR B0656+14 with eROSITA

### v1

@@ Abstract
Extended very-high-energy $\gamma$-ray emission from middle-aged pulsars as revealed recently by several groundbased $\gamma$-ray experiments has strong implication on the transport of high-energy particles in the interstellar medium surrounding those pulsars. The $\gamma$-ray emission is widely believed to be produced by high-energy electrons and positrons accelerated by the pulsar wind nebulae when scattering off the interstellar radiation field via the inverse Compton process. Multiwavelength counterparts of the $\gamma$-ray halos are expected to be present, which are, however, not observed yet. In this work, we report for the first time the detection of extended X-ray emission from $\sim 0.2^{\circ}$ radius region of PSR B0656+14 with eROSITA. The spectrum of the emission can be described by a power-law function with an index of $\sim 3.7$. The fluxes decrease with radius faster than the prediction of the particle diffusion and synchrotron radiation in a uniform magnetic field, suggesting the existence of a radial gradient of the magnetic field strength as $\sim r^{-1}$. The magnetic field strength in the X-ray emitting region is constrained to be $4-10~{}\mu$G.
1 Introduction
@@
Identifying multi-wavelength counterparts of the VHE pulsar halos are particularly important in understanding the emission mechanism of the halos and constraining the magnetic properties of the surrounding ISM. Possible extended GeV $\gamma$-ray emitting halo around Geminga in the Fermi-LAT data has been reported (Di Mauro et al., 2019). However, a controversial result was given in Xi et al. (2019), making this detection inconclusive. Efforts to search for X-ray and radio counterparts have been paid, but no clear detection has been found yet (Liu et al., 2019a; Khokhriakova et al., 2024; Manconi et al., 2024; Wu et al., 2024; Hooper et al., 2024).
In this work, we report the detection of an extended X-ray halo in a region with radius of $0.2^{\circ}$ around the Monogem pulsar (PSR B0656+14, Right Ascension $=104.951117^{\circ}$, Declination $=+14.239175^{\circ}$), with the eROSITA data taken during its Performance Verification (PV) phase (Schwope et al., 2022). The PWN associated with PSR B0656+14 as revealed by Chandra is about $15^{\prime\prime}$ (Bîrzan et al., 2016), which is much smaller than the emission size found in this work. The most natural explanation of the extended X-ray halo is the synchrotron radiation of electrons and positrons which give rise to the VHE $\gamma$-ray emission. The magnetic field strength is thus constrained to be $4-10~{}\mu$G based on the multi-wavelength data.
2 Data Reduction
@@
The eROSITA observations (OBS_ID: 300000) of PSR B0656+14 were performed on October 14th, 2019 in pointing mode [^1] We didn’t include the data in survey mode from the eROSITA Data Release 1(DR1) due to very limited exposure time and quite different instrumental response. during Calibration and Performance Verification (Cal-PV) phase. Due to the operation, telescope module (TM) 1 failed to function properly, while TM 5 and TM7 were affected by optical light leaking, as they do not equip an aluminium optical light filter (Predehl et al., 2021). To ensure optimal sensitivity in the soft X-ray band and extend the low-energy range down to 0.2 keV, we utilized data only from TM2, TM3, TM4, and TM6, i.e., TM8 products. The data were processed using the eROSITA Science Analysis Software System (eSASS) for DR1. Periods of high background activity were excluded by employing the flaregti task, yielding an effective exposure of 99.99 ks.
@@
For spectral extraction, we defined an annular region centered on PSR B0656+14 with radii from 4 to 26 to avoid central source contamination and minimize uncertainties due to edge effects of the instrument’s field of view. Within this region, approximately 530 point sources were identified and excluded using the ermldet tool (Fig. A1; see Appendix A for more details). This annular region was subsequently divided into 12 concentric rings, each with a radius increment of 2, resulting in 12 spectra and their corresponding response files. When generating the response matrices and auxiliary response files, the parameter exttype and psftype of the tool srctool were specifically set to “tophat” and “none” with assuming the the diffuse emission region are very extended. Each spectrum includes an instrumental background spectrum derived from filter wheel closed data. Ultimately, we obtained 12 spectra spanning an energy range of 0.2–2 keV. These spectra are re-binned to ensure sufficient photon counts per channel for better statistical reliability (see Fig. A2 in Appendix A).
3 Spectral fitting and results
@@
The diffuse X-ray background is theoretically a composite of several components (e.g., Ponti et al., 2023). The first is the cosmic X-ray background (CXB) predominantly originating from active galactic nuclei and galaxy clusters. Its spectrum is characterized by a power-law distribution with a fixed slope of $\sim 1.46$ and a fixed normalization of $\sim 11.6\,\rm photons$ $\rm cm^{-2}s^{-1}sr^{-1}keV^{-1}$ at 1 keV. Its foreground absorption is modeled based on the HI column density of the Milky Way. The second background is the emission from the local hot bubble (LHB) and the solar wind charge exchange (SWCX). In CCD spectra, these components can be modeled using an APEC plasma model with a temperature of $\sim$0.1 keV, assuming solar metallicity and no foreground absorption (Ponti et al., 2023). The third one is the emission from the Milky Way’s circum-galactic medium (CGM), Galactic halo, and unresolved stellar contributions. The CGM contains a hot, tenuous plasma, with a characteristic temperature of 0.15-0.2 keV and low metallicity. Some high-temperature plasma around 0.4-0.7 keV in the halo or the Galactic corona and unresolved M-dwarf stars contribute soft X-rays in the 0.6-2 keV band. Consequently, an empirical fitting approach is employed, utilizing two absorbed APEC components for the third background emission: one with an initial temperature around 0.2 keV with free metallicity and the other is fixed at 0.7 keV with solar abundance. We fix the temperature and metallicity of the latter one due to the insufficient constraints of the data as Yeung et al. (2024). Both APEC components (CGM and Galactic halo) and CXB share the same fixed absorption HI column density of $1.28\times 10^{21}\,\rm cm^{-2}$ from the HI observation. In this work, we use the tbabs model (Wilms et al., 2000) to estimate the ISM absorption. Finally, the adopted model of the background components is $\tt(tbabs*(powerlaw_{CXB}+apec_{CGM}+apec_{Halo})+apec_{LHB})$.
@@
To further verify the detected excess component in the inner three annuli, we use the spectrum of the $13^{\prime}-16^{\prime}$ annulus as data-driven background and re-fit the spectra of the inner three annuli. In this procedure, the instrumental background (i.e. filter wheel closed (FWC) data [^3] https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/FWC_dr1) has been subtracted individually for both the signal and background regions. In addition, the BACKSCAL keywords of the three spectra are adjusted to the BACKSCAL columns to account for the energy-dependent differences in their ARFs relative to the background spectrum. The fitting results of model ${\tt tbabs_{diffuse}*powerlaw_{diffuse}}$ for these three annuli are shown in Fig. 2 and summarized in Table 1. The results are consistent with the joint fitting of the 12 spectra within uncertainties, further supporting the detection of an excess power-law component in the $4^{\prime}-10^{\prime}$ region around PSR B0656+14.
The central pulsar is very bright in X rays. Due to the limited point spread function (PSF) of the instrument, the central source may contaminate the emission in outer regions. We estimate that the contribution from the central pulsar to the detected power-law excess due to the PSF can be safely ignored in the $4^{\prime}-6^{\prime}$ region (with possible contamination being smaller than $10\%$; see Appendix A). Therefore this component is not due to the contamination from the central point source. The results in Fig. 1 suggest the presence of an additional diffuse component in the inner three annuli. While such diffuse emission may exist in the outer annuli, its detection is beyond the sensitivity of the current observation.
@@
$4-6$ $2.08^{+1.36}_{-1.16}$ $3.69^{+0.40}_{-0.36}$ $11.20^{+4.06}_{-2.67}$ 35.93/54
$6-8$ $3.64^{+1.47}_{-1.01}$
$8-10$ $2.05^{+1.01}_{-0.73}$
Notes: Errors are at 90% confidence level. Columns from left to right are: annulus region, HI column density, photon power-law index, flux in 0.2 - 2.0 keV, and fitting $\chi^{2}$ value divided by the number of degree-of-freedom (dof).
@@
Using the data from eRASS:4 (stacked data from the first four consecutive all-sky surveys), Khokhriakova et al. (2024) searched for extended X-ray emission from 5 middle-aged pulsars, and found no significant diffuse emission. The exposure in their analysis is about 500 s for PSR B0656+14, which is perhaps too short to give a sensitive detection. The flux upper limit obtained is about $1.89\times 10^{-12}$ erg s-1 cm-2 deg-2 in the $0.5^{\circ}-1.0^{\circ}$ region and 0.5 - 2.0 keV band, assuming $\Gamma=2$. Their region of interest is beyond the signal region in this work, and the comparison is not direct. Furthermore, the sensitivity of point source detection is not as high as in this analysis, and the unresolved source population may result in a too high background estimate, which affects the detection of possible excess. Using the XMM-Newton observations, an upper limit of $3.3\times 10^{-14}$ erg s-1 cm-2 in the 0.5 - 5.0 keV band within $400^{\prime\prime}$ to $750^{\prime\prime}$ annulus was derived (Wu et al., 2024). In the same sky region and energy band, we obtain a flux of $2.70^{+0.91}_{-0.69}\times 10^{-14}$ erg s-1 cm-2, which is consistent with the XMM-Newton limit.
4 Interpretation
@@
High-energy electrons and positrons are accelerated by the PWN, which are then injected into the ISM. These particles propagate in the surrounding ISM, and interact with the background radiation field and magnetic field, producing gamma-ray emission via the inverse Compton scattering (ICS) process and X-ray emission via the synchrotron process. The injection spectrum is assumed to be a super-exponentially cutoff power-law function, $Q(E_{e})=Q_{0}E_{e}^{-\alpha}\exp[-(E_{e}/E_{\rm cut})^{\beta}]$, and the injection rate is assumed to be proportional to the spin-down luminosity of the pulsar, $L_{\rm sd}(t)=L_{0}(1+t/\tau)^{-2}$, where $\tau\approx 10^{4}$ yr is a characteristic time scale of the evolution of the pulsar spin-down luminosity. The diffusion coefficient is assumed to be spatially uniform, energy-dependent, as $D(E_{e})=D_{0}(E_{e}/{\rm GeV})^{\delta}$. Here we do not introduce the non-uniform diffusion (Fang et al., 2018) or anisotropic diffusion (Liu et al., 2019b) since we focus on relatively small spatial scales (within $10^{\circ}$ around the pulsar). The diffusion coefficient is adopted to be $D(100~{}{\rm TeV})=4.5\times 10^{27}$ cm2 s-1 as inferred from VHE gamma-ray observations (Abeysekara et al., 2017), and $\delta=1/3$ is assumed. Electrons and positrons experience strong energy losses during the propagation. We include the synchrotron cooling and the ICS cooling with three gray-body components of the background radiation field (Abeysekara et al., 2017): the cosmic microwave background with temperature $T=2.73$ K and energy density $u=0.26$ eV cm-3, an infrared background with $T=20$ K and $u=0.3$ eV cm-3, and an optical background with $T=5000$ K and $u=0.3$ eV cm-3. The ICS cooling rate is calculated following the new formalism of Fang et al. (2021).
@@
The model predicts a relatively flat radial profile, if a constant magnetic field strength ($B=4.2~{}\mu$G) is assumed, as shown by the dashed in the right panel of Fig. 3. The eROSITA data shows a fast decrease trend of the emission, different from the model prediction. It may indicate that there is a radial gradient of the magnetic field. Assuming that [^4] Here we assume that the magnetic field approaches to a constant ISM value of 2.5 $\mu$G at large distance from the pulsar. $B(r)={\rm max}[3.3(r/{\rm pc})^{-1},2.5]~{}\mu$G, the measured radial profile can be reproduced, as shown by the solid line in Fig. 3.
5 Conclusion and discussion
@@
The spectrum of the X-ray excess is somehow soft compared with the HAWC gamma-ray observations. The current simple isotropic diffusion model seems difficult to explain both the X-ray and gamma-ray spectra well. Improved measurements of the gamma-ray spectrum with higher sensitivity and better energy resolution by e.g., LHAASO (Ma et al., 2022) can better clarify the spectral behavior at VHE band. From the modelling point of view, as shown in Wu et al. (2024), the anisotropic diffusion model may give softer X-ray spectrum than the isotropic model while keeping similar TeV gamma-ray spectra.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (Nos. 12220101003, 12321003, 12273112), the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061).

### v2

# Possible evidence for extended X-ray emission surrounding PSR B0656+14 with eROSITA

@@ Abstract
[…]
Cosmic rays — Gamma-ray observatories — High energy astrophysics — Non-thermal radiation sources — Pulsars
I Introduction
@@
Identifying multi-wavelength counterparts of the VHE pulsar halos is particularly important in understanding the emission mechanism of the halos and constraining the magnetic properties of the surrounding ISM. Possible extended GeV $\gamma$-ray emitting halo around Geminga in the Fermi-LAT data has been reported (Di Mauro et al., 2019), which is, however, subject to debate (Xi et al., 2019). Efforts to search for X-ray and radio counterparts have been paid, but no clear detection has been found yet (Liu et al., 2019a; Khokhriakova et al., 2024; Manconi et al., 2024; Wu et al., 2024; Hooper et al., 2024)  [^1] The extended $\gamma$-ray source HESS J1809-193 was recently suggested to be a candidate pulsar halo (H. E. S. S. Collaboration et al., 2023b) with an X-ray halo (Li et al., 2023). However, this region is very complicated, and the origin of the $\gamma$-ray emission is ambiguous..
We note that the spectral similarity between the central pulsar and the diffuse emission may result from model degeneracies, since weak components can often be fitted acceptably by multiple models. The leakage fraction beyond $4^{\prime}$ is expected to vary with photon energy and should not reproduce the pulsar spectrum. Moreover, the normalization values, serving as a proxy for the source flux, differ significantly between the pulsar model and the diffuse emission(Khokhriakova et al., 2025), further suggesting that the diffuse component is not dominated by PSF leakage.
[…]
II Data Reduction
@@
The eROSITA observations (OBS_ID: 300000) of PSR B0656+14 were performed on October 14th, 2019 in pointing mode [^2] We did not include the data in survey mode from the eROSITA Data Release 1 (DR1) due to very limited exposure time and quite different instrumental response. during Calibration and Performance Verification (Cal-PV) phase. Due to the operation, telescope module (TM) 1 failed to function properly, while TM 5 and TM7 were affected by optical light leaking, as they do not equip an aluminium optical light filter (Predehl et al., 2021). To ensure optimal sensitivity in the soft X-ray band and extend the low-energy range down to 0.2 keV, we utilized data only from TM2, TM3, TM4, and TM6. The data were processed using the eROSITA Science Analysis Software System (eSASS) for DR1. Periods of high background activity were excluded by employing the flaregti task, yielding an effective exposure of 99.99 ks.
@@
For spectral extraction, we defined an annular region centered on PSR B0656+14 with radii from 4 to 26 to avoid central source contamination and minimize uncertainties due to edge effects of the instrument’s field of view. Within this region, approximately 530 point sources were identified and excluded using the ermldet tool (Fig. A1; see Appendix A for more details). This annular region was subsequently divided into 12 concentric rings, each with a radius increment of 2, resulting in 12 spectra and their corresponding response files. When generating the response matrices and auxiliary response files, the parameter exttype and psftype of the tool srctool were specifically set to “tophat” and “none” with assuming the the diffuse emission region are very extended. Each spectrum includes an instrumental background spectrum derived from filter wheel closed (FWC) data. Ultimately, we obtained 12 spectra spanning an energy range of 0.2–2 keV. These spectra are re-binned to ensure sufficient photon counts per channel for better statistical reliability (see Fig. A2 in Appendix A).
III Spectral fitting and results
@@
The diffuse X-ray background is theoretically a composite of several components (e.g., Ponti et al., 2023). The first is the cosmic X-ray background (CXB) predominantly originating from active galactic nuclei and galaxy clusters. Its spectrum is characterized by a power-law distribution with a fixed slope of $\sim 1.46$ and a fixed normalization of $\sim 11.6\,\rm photons$ $\rm cm^{-2}~s^{-1}~sr^{-1}~keV^{-1}$ at 1 keV. Its foreground absorption is modeled based on the HI column density of the Milky Way. The second background is the emission from the local hot bubble (LHB) and the solar wind charge exchange (SWCX). In CCD spectra, these components can be modeled using an APEC plasma model with a temperature of $\sim$0.1 keV, assuming solar metallicity and no foreground absorption (Ponti et al., 2023). The third one is the emission from the Milky Way’s circum-galactic medium (CGM), Galactic halo, and unresolved stellar contributions. The CGM contains a hot, tenuous plasma, with a characteristic temperature of 0.15-0.2 keV and low metallicity. Some high-temperature plasma around 0.4-0.7 keV in the halo or the Galactic corona and unresolved M-dwarf stars contribute soft X-rays in the 0.6-2 keV band.
Consequently, an empirical fitting approach is employed, utilizing two absorbed APEC components for the third background emission: one with an initial temperature around 0.2 keV with free metallicity and the other is fixed at 0.7 keV with solar abundance. We fix the temperature and metallicity of the latter one due to the insufficient constraints of the data as Yeung et al. (2024). Both APEC components (CGM and Galactic halo) and CXB share the same fixed absorption HI column density of $1.28\times 10^{21}\,\rm cm^{-2}$ from the HI observation. In this work, we use the tbabs model (Wilms et al., 2000) to estimate the ISM absorption. Finally, the adopted model of the background components is $\tt(tbabs*(powerlaw_{CXB}+apec_{CGM}+apec_{Halo})+apec_{LHB})$.
@@
To further verify the detected excess component in the inner three annuli, we use the spectrum of the $13^{\prime}-16^{\prime}$ annulus as data-driven background and re-fit the spectra of the inner three annuli. In this procedure, the instrumental background, i.e. FWC data [^4] https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/FWC_dr1/, has been subtracted individually for both the signal and background regions. In addition, the BACKSCAL keywords of the three spectra are adjusted to the BACKSCAL columns to account for the energy-dependent differences in their ARFs relative to the background spectrum. The fitting results of model $\tt tbabs_{diffuse}*powerlaw_{diffuse}$} for these three annuli are shown in Fig. 2 and summarized in Table 1. The results are consistent with the joint fitting of the 12 spectra within uncertainties, further supporting the detection of an excess power-law component in the $4^{\prime}-10^{\prime}$ region around PSR B0656+14.
The central pulsar is very bright in X-rays. Due to the limited point spread function (PSF) of the instrument, the central source may contaminate the emission in outer regions. The PSF model based on groundbased calibration (Brunner et al., 2022) is limited in accuracy at large radii. We therefore choose a pointing-mode observation of a bright variable star, V* V702 CrA, to test the effect of the PSF leakage. Note that, V* V702 CrA was observed at an off-axis angle of $4.5^{\prime}$, which is even bigger than the off-axis angle of PSR B0656+14 ($2.5^{\prime}$). Instrumentally, a larger off-axis angle is expected to give an even broader PSF. As shown in Fig. A4 of Appendix B, the radial profile of V* V702 CrA is more concentrated than that of PSR B0656+14 in the 0.2–2.0 keV band, supporting the presence of the diffuse emission. The fractional energy beyond $4^{\prime}$ of V* V702 CrA agrees well with the PSF model prediction. We estimate that the contribution from the central pulsar to the detected power-law excess can be neglected in the $4^{\prime}-6^{\prime}$ region, with possible contamination being smaller than 10% given the excess flux is about 1.4% of the central pulsar’s flux (see Appendix A). As a comparison, an estimated fraction of $\sim 2.7\%$ of the central source is expected beyond $4^{\prime}$, according to the eROSITA survey-mode PSF (Churazov et al., 2023). This indicates that the PSF in the survey mode is noticeably degraded due to the instrumental blurring compared to that in the pointing mode around the optical axis. These results indicate that the observed extended excess around PSR B0656+14 is not likely to be an artifact of the PSF. While a well-calibrated PSF remains crucial for such analyses, future deeper observations and improved characterization of the eROSITA pointing PSF will help confirm this conclusion.
@@
464-6 $2.08^{+1.36}_{-1.16}$ $3.69^{+0.40}_{-0.36}$ $11.20^{+4.06}_{-2.67}$ 35.93/54
686-8 $3.64^{+1.47}_{-1.01}$
8108-10 $2.05^{+1.01}_{-0.73}$
@@
Using the data from eRASS:4 (stacked data from the first four consecutive all-sky surveys), Khokhriakova et al. (2024) searched for extended X-ray emission from 5 middle-aged pulsars, and found no significant diffuse emission. The exposure in their analysis is about 500 s for PSR B0656+14, which is perhaps too short to give a sensitive detection. The flux upper limit obtained is about $1.89\times 10^{-12}$ erg s-1 cm-2 deg-2 in the $0.5\arcdeg-1.0\arcdeg$ region and 0.5 - 2.0 keV band, assuming $\Gamma=2$. Their region of interest is beyond the signal region in this work, and the comparison is not direct. Furthermore, the sensitivity of point source detection is not as high as in this analysis, and the unresolved source population may result in a too high background estimate, which affects the detection of possible excess. Using the XMM-Newton observations, an upper limit of $3.3\times 10^{-14}$ erg s-1 cm-2 in the 0.5 - 5.0 keV band within $400^{\prime\prime}$ to $750^{\prime\prime}$ annulus was derived (Wu et al., 2024). In the same sky region and energy band, we obtain a flux of $2.70^{+0.91}_{-0.69}\times 10^{-14}$ erg s-1 cm-2, which is consistent with the XMM-Newton limit. These results confirm the robustness of the detected diffuse component and motivate further modeling of its physical origin, as discussed below.
IV Interpretation
@@
High-energy electrons and positrons are accelerated by the PWN, which are then injected into the ISM. These particles propagate in the surrounding ISM, and interact with the background radiation field and magnetic field, producing gamma-ray emission via the inverse Compton scattering (ICS) process and X-ray emission via the synchrotron process. The injection spectrum is assumed to be a super-exponentially cutoff power-law function, $Q(E_{e})=Q_{0}E_{e}^{-\alpha}\exp[-(E_{e}/E_{\rm cut})^{\beta}]$, and the injection rate is assumed to be proportional to the spin-down luminosity of the pulsar, $L_{\rm sd}(t)=L_{0}(1+t/\tau)^{-2}$, where $\tau\approx 10^{4}$ yr is a characteristic time scale of the evolution of the pulsar spin-down luminosity. The diffusion coefficient is assumed to be spatially uniform, energy-dependent, as $D(E_{e})=D_{0}(E_{e}/{\rm GeV})^{\delta}$. Here we do not introduce the non-uniform diffusion (Fang et al., 2018) or anisotropic diffusion (Liu et al., 2019b) since we focus on relatively small spatial scales (within 10° around the pulsar). The diffusion coefficient is adopted to be $D(100~{\rm TeV})=4.5\times 10^{27}$ cm2 s-1 as inferred from VHE gamma-ray observations (Abeysekara et al., 2017), and $\delta=1/3$ is assumed. Electrons and positrons experience strong energy losses during the propagation. We include the synchrotron cooling and the ICS cooling with three gray-body components of the background radiation field (Abeysekara et al., 2017): the cosmic microwave background with temperature $T=2.73$ K and energy density $u=0.26$ eV cm-3, an infrared background with $T=20$ K and $u=0.3$ eV cm-3, and an optical background with $T=5000$ K and $u=0.3$ eV cm-3. The ICS cooling rate is calculated following the new formalism of Fang et al. (2021).
@@
[…]
V Conclusion and discussion
@@
The identification of weak extended emission is subject to uncertainties of the PSF model due to the lack of on-orbit data of the pointing observation mode. A better understanding of the on-orbit PSF is very important in confirming the results reported in this work. Physically the extended X-ray emission associated with ultra-high energy $\gamma$-ray sources is not unexpected. For instance, Suzuki et al. (2025) have recently reported the detection of X-ray diffuse emission around the PeVatron microquasar V4641 Sgr with XRISM observation, supporting the possibility of extended high-energy halos associated with different types of compact sources. Multi-wavelength observations are important in pinning down the radiation nature of these source.
The spectrum of the X-ray excess appears softer compared with the HAWC $\gamma$-ray observations. The simple isotropic diffusion model thus has difficulty in simultaneously explaining both the X-ray and $\gamma$-ray spectra. Improved measurements of the $\gamma$-ray spectrum with higher sensitivity and better energy resolution, e.g., by LHAASO (Ma et al., 2022), will help clarify the spectral behavior at the VHE band. From the modelling perspective, as shown in Wu et al. (2024), the anisotropic diffusion scenario may produce a softer X-ray spectrum than the isotropic case while keeping similar TeV $\gamma$-ray spectra. Future high-resolution X-ray missions such as Athena will also provide crucial constraints on the diffuse emission.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (Nos. 12220101003, 12321003, 12273112) and the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061).